OpenStudy (anonymous):

Help me through this! Giving medal! What is the equation in point-slope form of the line passing through (–1, 3) and (1, 7)?

OpenStudy (studygurl14):

First, use the slope formula: \(\huge \frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (studygurl14):

With points on the line \(\large (x_1,y_1)\) and \(\large (x_2, y_2)\)

OpenStudy (anonymous):

7-3 --- 1-(-1) 4/2 Slope is 2

OpenStudy (studygurl14):

nice job. Now plug what you know into the point-slope formula for the equation of a line \(\large y-y_1=m(x-x_1)\) where m = slope

OpenStudy (anonymous):

y-7 = 2(x-1)

OpenStudy (anonymous):

So is that the answer?

OpenStudy (triciaal):

yes point slope format of the equation of a line